# Efficient tensor network simulation of IBM’s largest quantum processors

### v1

@@ Abstract
Acknowledgements.- We acknowledge Donostia International Physics Center (DIPC), Ikerbasque, Basque Government, Diputación de Gipuzkoa and European Innovation Council (EIC) for constant support, as well as insightful discussions with the technical teams from Multiverse Computing and DIPC on the algorithms and technical implementations. We also acknowledge Miles Stoudenmire and Joseph Tindall for providing access to previous data. S. S. J. also acknowledges the Institute for Advanced Studies in Basic Sciences (IASBS).
Supplementary material: Weight-$N$ observables
@@
In the above equations, integers $P(size)$ and $Q(size)$ represent qubit labels depending on the size, as shown in the table below.
@@
Vertices are labeled sequentially from left to right and top to bottom, starting with 0.

### v3

@@ VI Acknowledgments
We acknowledge DIPC, Ikerbasque, Basque Government, Diputación de Gipuzkoa and EIC for support, as well as discussions with the teams from Multiverse Computing and DIPC. We also acknowledge Miles Stoudenmire and Joseph Tindall for providing access to previous data, and also comments by the community. S. S. J. also acknowledges IASBS. Simulations in Figs.(3(b),5,6) were done in the ATLAS cluster.
Appendix A Results near critical point
In the range of $\theta_{h}\in(\pi/8,\pi/4)$ depicted in Fig 2(b), our gPEPS results with lower bond dimension are more accurate than the higher ones, contrary to the monotonic improvement of accuracy with the increament of bond dimension. The underlying reason is the presence of a critical point ($\theta_{h}\approx 0.6$) in the vicinity, where the correlation length diverges. Our simple update-based gPEPS approach, along with local measurements, struggles to capture the correlations adequately with smaller bond dimensions, leading to overestimation or underestimation. Nonetheless, we have confirmed the convergence of our observable expectation values towards higher bond dimension.
Appendix B Weight-$N$ observables
@@
In the above equations, integers $P(size)$ and $Q(size)$ represent qubit labels depending on the size, as shown in the table below. Vertices are labeled sequentially from left to right and top to bottom, starting with 0.
@@
Appendix C Local tree structure of heavy-hexagon lattice
Figure 8: (Color online) Comparing a patch of heavy-hexagon lattice (left) with square lattice (right).
Simple update and local measurements are most effective for lattices exhibiting a local tree structure and shorter-range correlations, owing to the absence of loops. In the Figs.8, we present a comparison between a patch of 2D heavy-hexagon lattice and a 2D square lattice, illustrating the environment of the blue edge (pair of spins). It is evident that for a 2D square lattice, the ‘n-th’-neighbor environment contains loops for $n>1$. Conversely, in the case of a heavy-hexagonal lattice, the ‘n-th’-neighbor environment exhibits loops for $n>5$. Thus, for models with short-range correlations, such as the Ising transverse field away from the critical point, the heavy-hexagon lattice behaves akin to a tree structure locally. Consequently, local tensor updates employed in gPEPS, as well as local measurements, excel in capturing the real-time dynamics of the kicked Ising experiment.